8. 2/6 points I Previous Answers SwokATG13 4.3.018 Find the zeros of f(x), and state the multiplicity of each zero. (Order your answers from smallest to largest x.) fx) x(x 4)(2x 7) X with multiplicity 1 with multiplicity 3 x = 1-4 with multiplicity 2 Need Help? Read ItWatchtTalk to a Tutor Submit Answer Save Progress Practice Another Version -4 points SwokATG13 4.3,019 O Type here to search O e 99+
Answer #1

Solution:

\small f\left ( x \right )=x\left ( x+4 \right )^{3}\left ( 2x-7 \right )^{2}

Put   \small f\left ( x \right )=0

\small \therefore x\left ( x+4 \right )^{3}\left ( 2x-7 \right )^{2}=0

\small \therefore x=0

\small \left ( x+4 \right )^{3}=0\Rightarrow x=-4\: \: \: \left ( thrice \right )

\small \left ( 2x-7 \right )^{2}=0\Rightarrow x=\frac{7}{2},\frac{7}{2}

\small {\color{DarkRed} x=-4}  with multiplicity \small {\color{DarkRed} 3}

\small {\color{Magenta} x=0}   with multiplicity \small {\color{Magenta} 1}

\small {\color{DarkGreen} x=\frac{7}{2}} with multiplicity \small {\color{DarkGreen} 2}